### v1

# Why aphids are not pests in cacao? An approach based on a predator-prey model with aging

Abstract
We studied a mean-field predator-prey model with aging to simulate the interaction between aphids (Toxoptera aurantii) and syrphid larvae in cacao farms in Ilheus, Bahia. Based on the classical predator-prey model, we propose a system of differential equations with three rate equations. Unlike the original Lotka-Volterra model, our model includes two aphid population classes: juveniles (non-breeding) and adult females (asexually breeding). We obtained steady-state solutions for juvenile and adult populations by analyzing the stability of the fixed points as a function of model parameters. The results show that the absorbing state (zero prey population) is always possible, but not consistently stable. A nonzero stationary solution is achievable with appropriate parameter values. Using phase diagrams, we analyzed the stationary solution, providing a comprehensive understanding of the dynamics involved. Simulations on complete graphs yielded results closely matching the differential equations. We also performed simulations on random networks to highlight the influence of network topology on system behavior.
keywords:
aphids , cacao cultivation , population dynamics , mean-field approximation , differential equations
journal: Journal of Theoretical Biology
[…]
[dfisuefs] organization=Departamento de Física, Universidade Estadual de Feira de Santana,addressline=Av. Transnordestina S/N, Novo Horizonte, city=Feira de Santana, postcode=44031-460, state=Bahia, country=Brazil
[…]
[dcbiouefs] organization=Departamento de Ciências Biológicas, Universidade Estadual de Feira de Santana,addressline=Av. Transnordestina S/N, Novo Horizonte, city=Feira de Santana, postcode=44031-460, state=Bahia, country=Brazil
[…]
[dcnmeufscar] organization=Departamento de Ciências da Natureza, Matemática e Educação, Universidade Federal de São Carlos,addressline=Rodovia Anhanguera, Km 174, city=Araras, postcode=13600-970, state=São Paulo, country=Brazil
[…]
[ppgmuefs] organization=PPGM, Universidade Estadual de Feira de Santana,addressline=Av. Transnordestina S/N, Novo Horizonte, city=Feira de Santana, postcode=44031-460, state=Bahia, country=Brazil
1 Introduction
Studying predator-prey dynamics is essential for understanding ecological interactions and population control mechanisms in agricultural ecosystems [1, 2]. Various factors influence these dynamics, including environmental conditions, species interactions, and network structures [3]. In addition to these factors, the role of natural enemies in influencing the population dynamics of tropical insects has been well-documented [4, 5]. For instance, Godfray and Hassell (1987) argue that natural enemies may be a cause of discrete generations in tropical insects [6]. Moreover, recent studies have examined cacao plants’ interactions between aphid predators and ant-aphid mutualism [7, 8]. Silva and Perfecto (2013) explore how these interactions contribute to the coexistence of aphid predators, suggesting that ant-aphid mutualism may play a significant role in shaping predator-prey dynamics [9].
The relevance of classical predator-prey models to real-world scenarios has been questioned [10]. Dostálková, Kindlmann, and Dixon (2002) discuss the applicability of these models, highlighting their limitations and the need for incorporating more realistic ecological interactions [11]. Farina and Dennunzio (2008) extend this discussion by developing a predator-prey cellular automaton that includes parasitic interactions and environmental effects, providing a more nuanced understanding of these dynamics [12]. Furthermore, Stewart, Logsdon, and Kelley (2005) conducted an empirical study on the evolution of virulence under both horizontal and vertical transmission, providing insights into how these transmission modes influence pathogen dynamics and virulence [13]. Aphids are among the most serious pests, causing severe economic losses in several crops worldwide [14]. In cacao, however, the aphid Toxoptera aurantii is not perceived as a pest due to effective biological control by insect predators [15]. In cacao crops, the interactions between aphids and predators, mainly ladybugs, hoverflies, and lacewings, are poorly known [16].
One of the reasons why aphids cause low damage to cacao plants is the effectiveness of an assemblage of predators, primarily species of the families Coccinellidae (Coleoptera), Syrphidae (Diptera), Chrysopidae and Hemerobiidae (Neuroptera) [17]. In Ghana, Firempong and Kumar (1975) verified in field experiments that hoverflies and ladybug beetles were able to effectively suppress T. aurantii in cacao due to three factors: the synchronization of predator and prey colonization, the ability of predators to increase prey consumption at higher prey densities (functional response) and the complementary effect of the different coexisting species of natural enemies [18]. The effectiveness of biological control in various agricultural systems has been well-documented [19]. Carver (1978) highlighted the role of natural predators in controlling pest populations in several crops [20]. Numerous studies have highlighted the complex interactions within agricultural ecosystems [21, 22, 23]. Krebs (2001) provided a comprehensive overview of ecological methodologies and their applications in understanding these interactions [24].
The dynamics of natural regulation of insect pests through biological control is an important achievement for developing sustainable, ecologically based farming practices [25]. Modelling the dynamics of species interactions can be paramount to reaching such a level of understanding, as demonstrated for a natural-scale enemie interaction in shade coffee in Mexico [5]. Studying predator-prey interactions is fundamental to ecology, particularly in agroecosystems where pests can significantly impact crop yield and quality [26]. These environments are complex, with various species interacting, often resulting in intricate population dynamics [27]. These interactions can determine the success or failure of crop production, making it crucial to comprehend the underlying mechanisms that govern these relationships [23]. In many agricultural systems, certain insects can become pests, causing extensive crop damage [28]. For example, aphids are well-known pests in numerous agricultural contexts [29, 14, 30]. However, in cacao farms, the aphid Toxoptera aurantii is not perceived as a pest due to effective biological control by insect predators [29]. Aphids, specifically Toxoptera aurantii, are small sap-sucking insects that can reproduce rapidly, potentially leading to large infestations [31]. These infestations can reduce crop yields and quality, posing significant farmer challenges [32].
Aphids not only feed on plant sap, weakening the plant, but they also excrete honeydew, which can lead to the growth of sooty mold and attract other pests [14]. In many environments, the rapid reproduction and feeding behavior of aphids result in severe agricultural losses [33]. Their ability to transmit plant viruses further exacerbates their impact, making them one of the most formidable pests in agriculture [34]. However, intriguingly, in the cacao farms of Ilheus, Bahia, Brazil, aphids do not always manifest as severe pests [18, 9]. This observation suggests that there might be underlying ecological dynamics that prevent aphids from reaching pest status in these systems [35]. One possible explanation lies in the interactions between aphids and their natural predators, such as syrphid larvae (Syrphidae family). Syrphid larvae, commonly known as hoverfly larvae, are voracious predators of aphids [36]. They play a crucial role in biological control by preying on aphids and thus regulating their populations [37]. The effectiveness of syrphid larvae as biocontrol agents makes them significant in integrated pest management strategies [5]. Without predators, aphid population growth was typically logistic [38, 17, 39]. However, the voracious and early action of predators is responsible for pest population control before its typical exponential growth phase [37].
Through field experiments [9], differential equations (ODE) [40] and continuous-time computational simulations in this article, we study the effective control of aphids by hoverflies in cacao. To unravel this phenomenon, we developed a simple mean-field predator-prey model with aging dynamics [41, 42]. This model, proposed by our research team, simulates the intricate interplay between aphids and their natural predators in cacao ecosystems. Unlike traditional models such as the Lotka-Volterra framework, our study introduces another perspective by considering two distinct classes of aphid populations: juveniles, which are incapable of reproduction, and adult females, capable of asexual reproduction [43, 44]. It allows us to capture more accurately the life cycle and reproductive strategies of aphids, providing a more realistic depiction of the ecological dynamics [45, 46].
The main hypothesis of this study is that the presence of syrphid larvae effectively controls the aphid population, preventing them from becoming pests in cacao cultivation. Our research question is: How do the interactions between aphids and syrphid larvae influence the population dynamics of aphids in cacao farms, and why does this interaction prevent aphids from reaching pest status? Our approach involves the formulation of a system of differential equations with three rate equations, assuming a constant predator population. By analyzing steady-state solutions and the stability of fixed points, we aim to understand the conditions under which aphid populations are kept in check. Additionally, we perform simulations on both complete and random networks to study the phase transition from active to absorbing states and observe population dynamics differences. By delving into the intricacies of predator-prey interactions, this study aims to provide insights into why aphids may not always pose a threat as pests in cacao cultivation. Understanding these dynamics can inform more effective pest management strategies, contributing to sustainable agricultural practices.
2 Model
Since the pioneering studies of Volterra and Lotka [47, 48], the use of interacting populations in the prey-predator framework (also applicable to parasite-host interactions or consumption-resources) has generated a considerable amount of research aiming to understand the dynamics of this type of ecological interaction in different systems [6, 49]. Two main approaches have been applied in such studies: differential equations [50] and computational simulations [51].
Previous studies have demonstrated the importance of predator-prey dynamics in agricultural ecosystems. Firempong (1975 and 1976) and Kumar (1975) provided foundational insights into these interactions in cacao plantations [52, 18]. The mathematical modeling of ecological interactions has been extensively developed over the years. Turchin (2003) and Hassell (2000) offered significant contributions to the theoretical frameworks of population dynamics [53, 54]. In particular, Hassell’s work on the dynamics of arthropod predator-prey systems provides a comprehensive understanding of these interactions [55].
In this study, we developed a model for the dynamics of the interaction between the aphid T. aurantii and syrphid larvae, which were treated as effective biological control agents of this aphid. We propose a detailed mean-field predator-prey model incorporating aging dynamics to investigate the population dynamics between aphids and their natural predators in cacao farms. This model, based on the works of Nascimento Silva [9], simulates the interaction between aphids (Toxoptera aurantii) and syrphid larvae (Syrphidae), providing insights into why aphids do not reach pest status in these ecosystems.
We assume that the predator population, syrphid larvae, remains constant during the observation period. This assumption simplifies the model and focuses on the prey dynamics, which are influenced by predation and reproduction rates. The mean-field approximation is employed to reduce the complexity of the spatial interactions, assuming that each individual interacts with all others.
The key parameters of the model are:
    $\alpha$: the rate at which juveniles mature into adults, representing the maturation process of aphids;
    $\beta$: the predation rate on juveniles, indicating the effectiveness of predators on young aphids;
    $\gamma$: the reproduction rate of adult aphids, reflecting their asexual reproductive capabilities;
    $\epsilon$: the predation rate on adult aphids, showing the impact of predation on mature aphids;
    $k_{x}$: the number of neighboring sites, representing the local interaction range.
The model consists of a system of differential equations that describe the rate of change for juveniles $J_{x}(t)$, adults $A_{x}(t)$, and the available space $E_{x}(t)$ for the aphid population:
$\left\{\begin{array}[]{rcl}\displaystyle\frac{dJ_{x}(t)}{dt}&=&-\alpha J_{x}(t)-\beta J_{x}(t)+\frac{\gamma}{k_{x}}\sum_{y}P_{t}[E_{x}(t),A_{y}(t)]\,,\\ \displaystyle\frac{dA_{x}(t)}{dt}&=&\alpha J_{x}(t)-\epsilon A_{x}(t)\,,\\ \displaystyle\frac{dE_{x}(t)}{dt}&=&\beta J_{x}(t)-\frac{\gamma}{k_{x}}\sum_{y}P_{t}[E_{x}(t),A_{y}(t)]+\epsilon A_{x}(t)\,,\end{array}\right.$
where:
    $J_{x}(t)$ represents the probability of there being a juvenile aphid at the site $x$ at time instant $t$;
    $A_{x}(t)$ represents the probability of there being an adult aphid at the site $x$ at time instant $t$;
    $E_{x}(t)$ represents the probability that the site $x$ is empty and available for an aphid at time instant $t$.
    $P_{t}[E_{x}(t),A_{y}(t)]$ is the joint probability that, at time instant $t$, the site $x$ is empty and available for an aphid and there is an adult aphid at the site $y$.
We used numerical methods to solve the differential equations in the simple mean-field case, employing algorithms such as the fourth-order Runge–Kutta method implemented in the C programming language. The model was validated by comparing its predictions with empirical data from cacao farms and with results from other established models. Furthermore, we performed simulations on complete graphs and random networks, which are expected to yield results similar to those from simple mean-field calculations, thus testing the self-consistency of the model. Additionally, simulations on random networks explore the limitations of simple mean-field predictions by considering the inherent finite-size effects of the aphid population, providing a more realistic understanding of the model’s behavior.
3 Mean-field results
In this section, we employ a single-site mean-field approximation to simplify the complex interactions of our predator-prey model. Our goal is to analyze the population dynamics and system stability under various parameter settings. By focusing on average effects rather than detailed spatial dynamics, we aim to derive insights into the overall behavior of the model and identify conditions that prevent aphids from becoming pests. The mean-field approximation assumes that each individual interacts with the entire population, thus reducing the system’s complexity. This approach allows us to derive tractable analytical expressions and gain insights into the overall system behavior, which would be challenging to achieve with detailed spatial models.
We begin by setting $P_{t}[E_{x}(t),A_{y}(t)]\approx E_{x}(t)A_{y}(t)$ and assuming homogeneous populations of juveniles, adults, and empty spaces. Thus, we use $J_{x}(t)=J(t)$, $A_{x}(t)=A(t)$, $E_{x}(t)=E(t)$, and $A_{y}(t)=A(t)$ for all $x$ and $y$. With these assumptions, we approximate the interactions as follows:
$P_{t}[E(t),A(t)]$$\approx$ $E(t)A(t)\,,$
$\sum_{y}P_{t}[E_{x}(t),A_{y}(t)]$$\approx$ $k_{x}E(t)A(t)\,.$
This simplification leads to the following system of differential equations describing the rates of change for juveniles, adults, and empty spaces:
$\left\{\begin{array}[]{rcl}\displaystyle\frac{dJ(t)}{dt}&=&-\alpha J(t)-\beta J(t)+\gamma E(t)A(t)\,,\\ \displaystyle\frac{dA(t)}{dt}&=&\alpha J(t)-\epsilon A(t)\,,\\ \displaystyle\frac{dE(t)}{dt}&=&\beta J(t)-\gamma E(t)A(t)+\epsilon A(t)\,.\end{array}\right.$
3.1 Fixed points
Since solving this system of coupled differential equations analytically is complex, we instead find the fixed points by setting the derivatives to zero:
$\left\{\begin{array}[]{rcl}0&=&(-\alpha-\beta)J^{*}+\gamma[1-(J^{*}+A^{*})]A^{*}\,,\\ 0&=&\alpha J^{*}-\epsilon A^{*}\Rightarrow A^{*}=\alpha J^{*}/\epsilon.\end{array}\right.$
These equations yield two potential fixed points:
$\begin{array}[]{rcl}\left(J^{*}=0\right.&\textrm{and}&\left.A^{*}=0\right)\\ &\textrm{or}&\\ \left\{J^{*}=\displaystyle\frac{[\gamma\alpha-\epsilon(\alpha+\beta)]\epsilon}{\gamma\alpha(\epsilon+\alpha)}\right.&\textrm{and}&\left.A^{*}=\displaystyle\frac{\gamma\alpha-\epsilon(\alpha+\beta)}{\gamma(\epsilon+\alpha)}\right\}.\end{array}$
3.2 Stability analysis of fixed points
To determine the stability of these fixed points, we construct the Jacobian matrix $M$:
$M=\left[\begin{array}[]{cc}-\alpha-\beta-\gamma A^{*}&\gamma[1-(J^{*}+2A^{*})]\\ \alpha&-\epsilon\end{array}\right]\,.$
By calculating the eigenvalues of this matrix, we can assess the stability of each fixed point. The trivial fixed point $(J^{*}=0,A^{*}=0)$ represents the absorbing state where the aphid population is eradicated. This point loses stability when the components of the non-trivial fixed point become positive, indicating a transcritical bifurcation.
3.3 Bifurcation and phase diagrams
We observe that increasing the value of $\gamma$, the reproduction rate of adult aphids, leads to a transcritical bifurcation at a critical value $\gamma_{\textrm{c}}$, which depends on the other parameters $\alpha$, $\beta$, and $\epsilon$. The absorbing state (zero aphid population) remains a solution, but it loses stability as the system transitions to a state with a finite stationary population of aphids.
To visualize these transitions, we construct bifurcation and phase diagrams. The bifurcation diagram (Figure 1) shows the transition from the absorbing to a non-zero stationary state as $\gamma$ increases. The phase diagrams (Figures 2 and 3) in the $\beta\times\gamma$ and $\alpha\times\gamma$ planes illustrate the regions of parameter space where the system exhibits absorbing or active phases.
Figure 1: Bifurcation diagram showing the transition from the absorbing state (no aphids) to a non-zero stationary state (persistent aphid population). The parameters are $\alpha=1.5$, $\beta=1$, and $\epsilon=0.75$. Below are two examples of the evolution of the normalized aphid population for each of the phases: on the left, for the absorbing phase; on the right, for the active phase. The insets zoom into the transition regions, highlighting the critical value $\gamma_{\textrm{c}}$ where the shift occurs. These detailed views are essential to understand the precise behavior of the system at critical points.
In the bifurcation diagram, the x-axis represents the reproduction rate $\gamma$, and the y-axis shows the normalized aphid population. As $\gamma$ crosses the critical value $\gamma_{\textrm{c}}$, the system shifts from an absorbing phase (no aphids) to an active phase (persistent aphid population). This diagram provides a clear visualization of how the reproduction rate of adult aphids influences the overall population dynamics, indicating a critical threshold beyond which the aphid population becomes sustainable.
The phase diagrams further elucidate how changes in $\beta$ (predation rate on juveniles) and $\alpha$ (the rate at which juveniles mature) influence the system’s behavior. The results are presented in graphical form to illustrate the key findings of our analysis. Each graph includes insets to highlight specific aspects of the data that are crucial for understanding the overall results.
Figure 1 shows the bifurcation diagram with the x-axis representing the reproduction rate $\gamma$, and the y-axis showing the normalized aphid population. The insets in the graphs provide a zoomed-in view of critical regions that exhibit interesting or significant behavior, such as the transition from the absorbing state to a non-zero stationary state. These insets are necessary to understand the finer nuances of the model’s behavior and provide additional context to the main findings.
Figure 2 shows the phase diagram in the $\beta\times\gamma$ plane for fixed $\alpha=1.5$ and $\epsilon=0.75$. The diagram delineates the regions where the system is in an active (non-zero aphid population) or absorbing (zero aphid population) state. The insets illustrate the temporal evolution of the normalized aphid population corresponding to each phase, providing clear examples of how the populations behave over time in different parameter regimes.
Figure 3 illustrates the phase diagram in the $\alpha\times\gamma$ plane for fixed $\beta=1$ and $\epsilon=0.75$. It shows how variations in the rate at which juveniles become adults ($\alpha$) and the reproduction rate of adults ($\gamma$) affect the system’s state. The insets provide detailed views of the temporal evolution in different regions of the phase diagram, helping to understand the dynamic transitions between phases.
4 Simulation Results
We conducted extensive simulations on complete graphs and random networks to complement our mean-field analysis. These simulations verify theoretical predictions and offer nuanced insights into the system’s dynamics.
Figure 2: Phase diagram in the $\beta\times\gamma$ plane for fixed $\alpha=1.5$ and $\epsilon=0.75$. The diagram shows the regions where the system is in an active (non-zero aphid population) or absorbing (zero aphid population) state. In each inset, there is an example of the temporal evolution of the normalized aphid population corresponding to the phase where one is located in the diagram: on the top left, for the active phase; in the lower right, for the absorbing phase. The insets illustrate the different dynamic behaviors over time, clarifying the system’s response to changes in $\beta$ and $\gamma$.
4.1 Complete Graph Simulations
In simulations on complete graphs, every node is connected to every other node, which mirrors the mean-field assumption of uniform interactions. The results align well with the mean-field predictions, especially in the absorbing phase where the aphid population is zero. In the active phase, the simulated populations of juveniles and adults closely match the steady-state solutions predicted by the mean-field model. This consistency underscores the robustness of the mean-field approximation for large, well-mixed populations.
Figure 3: Phase diagram in the $\alpha\times\gamma$ plane for fixed $\beta=1$ and $\epsilon=0.75$. The diagram illustrates how variations in the rate at which juveniles become adults ($\alpha$) and the reproduction rate of adults ($\gamma$) affect the system’s state. In each inset, there is an example of the temporal evolution of the normalized aphid population corresponding to the phase where one is located in the diagram: on the top right, for the active phase; in the lower left, for the absorbing phase. The insets provide detailed temporal evolution, highlighting the critical thresholds that influence the system’s dynamics.
4.2 Random Network Simulations
Simulations on random networks, where connections between nodes are assigned randomly, show behaviors that complement the mean-field predictions. While the general trend of the phase transition from absorbing (no aphids) to active (persistent aphid population) states is consistent with the mean-field analysis, the exact parameter values at which these transitions occur can vary due to network-induced fluctuations and finite size effects. The transition points observed in simulations on random networks broadly align with those identified in the mean-field analysis. However, minor differences can occur due to the inherent variability in network structures.
Additionally, these simulations exhibit variability in exact population levels at equilibrium relative to mean-field predictions, highlighting the influence of network topology on the dynamics. Specifically, the population levels at equilibrium in the simulations do not exactly match those predicted by the mean-field model. There is noticeable variability in these equilibrium population levels, meaning that the equilibrium population can fluctuate or differ from the mean-field predictions due to factors such as network topology, random interactions, or other stochastic effects inherent in the simulations.
By systematically varying parameters $\alpha$, $\beta$, $\gamma$, and $\epsilon$ in the simulations, we mapped out the phase transitions in the system. The critical values at which the system transitions from the absorbing to the active phase are consistent with the bifurcation points identified in the mean-field analysis. However, the simulations demonstrate that random networks can sustain larger fluctuations.
4.3 Discussion
The simulation results provide valuable insights into the practical implications of predator-prey dynamics in agricultural settings. The ability of syrphid larvae to control aphid populations is confirmed, but the effectiveness of this biological control strategy can vary based on the spatial distribution and connectivity of the habitat. These findings suggest that enhancing habitat connectivity and promoting a well-mixed predator-prey environment could improve the stability and effectiveness of natural pest control mechanisms.
Overall, the simulation results reinforce the theoretical findings of the mean-field analysis while also uncovering additional layers of complexity introduced by network structure and finite-size effects. These insights are crucial for developing more nuanced and effective strategies for ecological management and pest control in agricultural ecosystems.
5 Conclusions
In this study, we investigated the predator-prey dynamics between aphids (Toxoptera aurantii) and syrphid larvae (Syrphidae) in cacao farms using a mean-field approximation and extensive simulations. Our goal was to understand why aphids do not reach pest status in these ecosystems and how natural predation controls their population.
Our mean-field model introduced a system of differential equations that incorporated key biological factors, such as the maturation process of aphids and their asexual reproduction. We derived steady-state solutions and analyzed the stability of the system’s fixed points. The results showed that the absorbing state, where the aphid population is zero, is a possible solution but not always stable. A nonzero stationary solution can be obtained with an appropriate set of parameters, as confirmed by the phase diagrams.
The simulations on complete graphs validated our mean-field predictions, demonstrating that the model accurately captures the system’s dynamics under well-mixed conditions. However, simulations on random networks revealed additional complexities. These simulations exhibited variability in the exact population levels at equilibrium relative to mean-field predictions, highlighting the influence of network topology on the dynamics. Specifically, random networks can sustain larger fluctuations, indicating that local interactions and stochastic effects play a significant role in population stability.
Our findings have practical implications for pest management in agricultural settings. The presence of syrphid larvae effectively controls aphid populations, preventing them from reaching pest status. The effectiveness of this biological control strategy may vary based on the spatial distribution and connectivity of the habitat. Enhancing habitat connectivity and promoting a well-mixed predator-prey environment could improve the stability and effectiveness of natural pest control mechanisms.
Overall, this study reinforces the theoretical foundations of predator-prey dynamics while uncovering the additional layers of complexity introduced by network structures. These insights are crucial for developing more nuanced and effective strategies for ecological management and pest control in agricultural ecosystems. By understanding the mechanisms that regulate aphid populations in cacao farms, we can better design sustainable agricultural practices that leverage natural predation for pest management [5].
Future research should explore the impact of different network topologies on predator-prey dynamics to better understand the role of spatial structure and connectivity in population stability. Empirical studies in real-world agricultural settings are essential to validate our theoretical and simulation results. Additionally, investigating interactions between multiple predator and prey species and incorporating environmental factors such as climate change and agricultural practices could provide a more comprehensive understanding of ecological dynamics [56]. Understanding the spatial dynamics of populations is crucial for effective pest management. Hanski (1999) and Levins (1969) have been pioneers in the study of metapopulations and their implications for ecology and conservation [57, 58].
Acknowledgements
V.R.V.A., N.G.F.M., E.N.S. and A.T.C.S. thank the Department of Education of the government of the State of Bahia for the incentive bonus for scientific production. E.N.S. thanks CAPES Foundation-Brazil for a scholarship and grants that allowed data collection in the field (BEX 1609/00-9). V.R.V.A. thanks Professor Dagoberto da Silva Freitas for encouraging him to return to scientific research. AC thanks the Centro de Ciências Agrárias of UFSCar for its support of scientific research.